OpenStudy (anonymous):

In 1800, the Michigan population of Mitchell's satyr was surveyed. There were 476 chestnut and 203 sandy colored butterflies. what is p and what is q

OpenStudy (anonymous):

Add the two together to get N = 476+ 203 = 679 p = proportion chestnut = 476/679 q = proportion sandy = 203/679